OpenStudy (anonymous):

Consider a spherical Gaussian surface of radius R centered at the origin. A charge Q is placed inside the sphere. Where should the charge be located to maximize the magnitude of the flux of the electric field through the Gaussian surface?

OpenStudy (anonymous):

*x=0 y=0 z=R/2 *x=0 y= R/2 z=0 *origin *x+R/2 yandz=0 *The charge can be located anywhere, since flux does not depend on the position of the charge as long as it is inside the sphere.

OpenStudy (anonymous):

Any help is appreciated

OpenStudy (anonymous):

Gauss's Law gives you the answer

OpenStudy (anonymous):

It states that: "The electric flux through any closed surface is proportional to the enclosed electric charge" What can you conclude?

OpenStudy (anonymous):

origin ?

OpenStudy (anonymous):

Does the law state any specific position or it is enough to have the charge enclosed?

OpenStudy (anonymous):

ohh okay so it doesnt matter where it is located? aslong as its enclosed

OpenStudy (anonymous):

If it is proportional to the charge and the charge is constant, then the flux is also constant, right? And if it is constant, then there is no maximum, it is always the same. Then, no matter where the charge is located, as far as it is enclosed, the flux will always be the same

OpenStudy (anonymous):

thank you so much...ive been rereading this chapter for an hour now ... thanks again

OpenStudy (anonymous):

Always focus on the physical concept. If the gaussian surface were not spherical, the result would be the same and if the radius were 100R instead of R, then again the same

OpenStudy (anonymous):

will doo..thanks again...im sure ill be back lol...Great help